# WildDESED: An LLM-Powered Dataset for Wild Domestic Environment Sound Event Detection System

### v1

@@ 1 Introduction
Building on this foundation, our research not only introduces the WildDESED dataset, but also explores the application of curriculum learning in the context of SED to tackle the challenges posed by domestic noisy environments. Curriculum learning [22, 23, 24] is a training approach that improves models for noisy speech [25, 26] and audio by starting with simpler, less noisy data and gradually increasing the noise level. This method is similar to the way how the humans learn and helps models adjust from clean to noisy sounds more effectively. In this work, we applied curriculum learning to the baseline convolutional recurrent neural network (CRNN) [27, 28, 29] model using the WildDESED dataset for our studies. The novelty of this work lies in the proposal of a new in-the-wild dataset for advancing SED research and exploring curriculum learning as an approach to develop noise-robust SED systems. We will publicly release the WildDESED dataset in due course.
@@ 3.3 How do we mix the background noises to the scenarios?
  • Sounds like ‘Footsteps’, ‘Door closing’, and ‘Chair moving’ fall into the Human-Related and Intermittent Sounds category. These are inserted at random intervals to simulate the sporadic nature of human movement and activities within a home. The volume and frequency of these sounds are varied to reflect the realistic and unpredictable nature of their occurrence in daily life.
@@
Figure 3 displays class-wise statistics for different background noises in each subset of the WildDESED dataset, indicating the prevalence of each noise type, within synth, synth val, weak, and test subsets. Figure 4 shows scenario-wise statistics for the scenarios in the WildDESED dataset, quantifying how frequently each scenario appears in each subset. Through this detailed dataset structure, WildDESED positions itself as a crucial resource for developing and evaluating SED systems, equipping researchers with the means to advance the field of SED in naturalistic home environments.

### v2

@@ 1 Introduction
Building on this foundation, our research not only introduces the WildDESED dataset, but also explores the application of curriculum learning in the context of SED to tackle the challenges posed by domestic noisy environments. Curriculum learning [22, 23, 24] is a training approach that improves models for noisy speech [25, 26, 27] and audio by starting with simpler, less noisy data and gradually increasing the noise level. This method is similar to the way how the humans learn and helps models adjust from clean to noisy sounds more effectively. In this work, we applied curriculum learning to the baseline convolutional recurrent neural network (CRNN) [28, 29, 30] model using the WildDESED dataset for our studies. The novelty of this work lies in the proposal of a new in-the-wild dataset for advancing SED research and exploring curriculum learning as an approach to develop noise-robust SED systems. The WildDESED dataset has been made publicly available [^1] https://github.com/swagshaw/WildDESED.
@@ 3.3 How do we mix the background noises to the scenarios?
  • Sounds like ‘Footsteps’, ‘Door closing’, and ‘Chair moving’ fall into the Human-Related and Intermittent Sounds category. These are inserted at random intervals to simulate the sporadic nature of human movement and activities within a home. The volume and frequency of these sounds are varied ±$\pm 10\%$ range to reflect the realistic and unpredictable nature of their occurrence in daily life.
@@
Figure 3 displays class-wise statistics for different background noises in each subset of the WildDESED dataset, indicating the prevalence of each noise type, within synth, synth val, weak, and test subsets. Figure 4 shows scenario-wise statistics for the scenarios in the WildDESED dataset, quantifying how frequently each scenario appears in each subset. Through this detailed dataset structure, WildDESED dataset positions itself as a crucial resource for developing and evaluating SED systems, equipping researchers with the means to advance the field of SED in diverse naturalistic home environments.